If an ounce of gold, valued at $760, increases at a rate of 8.1 percent per year, how long will it take to be valued at $1,000? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

Number of years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1000=760*(1.081)^n

(1000/760)=(1.081)^n

Taking log on both sides;

log (1000/760)=n*log (1.081)

n=log (1000/760)/log(1.081)

=3.52 years(Approx).